# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ III.1.1 Higgs Fields
where we take the trace in the fundamental representation. This is a lattice regularization of a gauged principal chiral model [33], a non-linear sigma model whose target space is the group manifold.
III.1.2 Hamiltonian Formulation and Gauss Law
@@
and $|\hat{E}_{\ell}|^{2}$ and $|\hat{Q}^{\text{m}}_{i}|^{2}$ are the corresponding quadratic Casimir operators, which do not depend on whether we use left- or right-generators. The operator $\hat{W}_{p}$ is the operator analog of Eq. 27, the trace of the oriented product of $\hat{U}_{\ell}$ on the links of spatial plaquette $p$. Similarly, $\hat{\Lambda}_{\ell}$ is the operator analog of Eq. 29.
The eigenstates of $|\hat{E}_{\ell}|^{2}$ correspond to the irreducible representations of the gauge group, with the $\hat{U}_{\ell}$ acting as raising and lowering operators [34, 35]. The same is true for the group-valued Higgs, with $\hat{\varphi}_{i}$ acting as the raising and lowering operators, while for the fundamental vector-valued Higgs, the charge eigenstates are angular momentum eigenstates of a rigid rotor.
@@
Note that this is similar but subtly distinct from the Abelian case, Eqs. 11, 12 and 13, where we used $\hat{E}_{\ell}^{\text{Abelian}}=-\hat{E}^{\text{Abelian}}_{-\ell}$. This relationship is not true in non-Abelian gauge theory. Instead, left- and right-translations of the gauge field are related by
@@
which are manifestly orientation-independent. The classic (though heuristic) way to think of this is that the gauge bosons (gluons) carry a distinct charge and anti-charge in the two directions along the link. In the Abelian case, $U_{\ell}^{\text{a}}=1$ in Eq. 39, and the link charge Eq. 40 is exactly zero.
@@
III.1.3 Open Boundaries and Global Symmetry
We introduce electric open boundary conditions as in Fig. 2, with dynamical links extending from the bulk to the vacuum which allow electric flux to pass through the boundary. We denote the link variables on the boundary link touching site $i$ by $U_{i}$, with the convention that the link is oriented “in” from the vacuum to site $i$. We have minimal open Wilson strings going around the boundary plaquettes, which we can write as
@@ III.1.3 Open Boundaries and Global Symmetry
Notice that each boundary link only receives a transformation from the inside end, where it terminates on a matter field. In addition to this gauge symmetry, the boundary action Eq. 42 has a physical global $\mathcal{G}$ symmetry acting on the boundary,
@@
which rotates all matter fields in the fundamental representation and all gauge fields, including boundary links, in the adjoint representation. But this is not the generator of global gauge transformations, Eq. 43, under which the boundary links only rotate from the inside. The generator of global gauge transformations is the “bulk charge”
@@
where we have assumed global gauge invariance to identify $\hat{Q}^{a}_{\text{bulk}}=0$. Thus the boundary symmetry Eq. 44 may be viewed, by the Gauss law constraint, as being generated by the total color charge of the system. This is analogous to the Abelian case, where the total charge of the system is just the matter charge, Eq. 15, since the links do not carry any charge. Note, however, that in the non-Abelian case the charge of the boundary links is “fractionalized” into a piece that contributes to the bulk charge and a piece that contributes to the boundary charge.
III.2 Boundary Symmetry Breaking in the SU(2) Higgs Model
@@
where $\phi_{1}$ and $\phi_{2}$ are complex numbers. Denoting the matrix Eq. 49 as $\varphi_{i}^{\text{f}}$, note that the determinant of this matrix is the length of the rotor. The Wilson link for the vector Higgs model, Eq. 28, can then be written as
$\Lambda_{ij}$ $=\langle\phi_{0},(\varphi^{\text{f}}_{i})^{-1}U_{ij}^{\text{f}}\varphi^{\text{f}}_{j}\phi_{0}\rangle$
@@
which is exactly equivalent to the group-valued Higgs definition, Eq. 29. This makes the SU(2) Higgs rotors special, since they may be viewed as either vector-valued or group-valued (which is also the case for U(1) rotor).
@@
III.2.1 Bulk Phase Diagram
To map the phase diagram, we perform classical Monte Carlo simulations for the 4D $\mathrm{SU}(2)$ model defined by the Fradkin-Shenker action, Eq. 7, with periodic boundary conditions on an $L^{4}$ hypercubic lattice. The phase diagram can be mapped out by measuring the Wilson plaquette and link observables along with their variances (i.e. susceptibilities), which are shown in Fig. 5. There is a roughly horizontal first-order transition line extending from a critical endpoint at $(\beta_{c},\kappa_{c})\approx(2.73,0.70)$ [36] (cyan box) towards $\beta\to\infty$. This line is clearly visible in the link susceptibility, Fig. 5(b).
The phase diagram exhibits only one clearly distinct bulk phase—the confined-Higgs phase—a deconfined phase being absent in non-Abelian theories. The phase diagram is, however, roughly separated into two regions indicated by the behavior of the Wilson link expectation value, shown in Fig. 5(a), with $\cramped{\langle\mathrm{Re}\,\Lambda_{\ell}\rangle\sim 0}$ indicating the confining regime and $\cramped{\langle\mathrm{Re}\,\Lambda_{\ell}\rangle\sim 1}$ indicating the Higgs regime. To the left of this critical endpoint is a supercritical region (a Widom line [38]) extending to smaller $\beta$ and larger $\kappa$, a rapid crossover from the Higgs to the confined regimes. The location of this supercritical region is most evident in the Wilson plaquette susceptibility, Fig. 5(d), which shows a pronounced intensity emanating from the critical endpoint. We expect this phase diagram to be qualitatively consistent with those for general non-Abelian Higgs models with either fundamental vector- or group-valued Higgs fields. In the group-valued case the continuity of the Higgs and confined regimes was proven by Fradkin and Shenker [12].
III.2.2 Boundary Symmetry Breaking
We now consider open boundaries, with boundary action Eq. 42 in a slab geometry. As in Section II.3, we start by considering the limiting behavior when $\kappa\to\infty$. In this limit every bulk link satisfies the constraint $\mathrm{Re}\,\Lambda_{\ell}=1$. In this section we will resolve this constraint by fixing a gauge. Gauge-invariant formulations for the group-valued representation are presented in Section III.3, and for the fundamental representation in Section III.4.
@@ III.2.2 Boundary Symmetry Breaking
[…]
@@
where the former is an $\mathrm{SU}(2)$ matrix which decomposes according to Eq. 52, and the latter is explicitly a 4-component unit-length vector. For large but finite $\kappa$, the bulk fluctuations are strongly gapped and the boundary should behave as quasi-$(D-1)$-dimensional, and we expect the boundary phase transition to persist as in the Abelian case.
@@
To test this prediction, we perform Monte Carlo simulations with open boundaries and measure the order parameter, Eq. 54 (defined in terms of gauge invariant observables Eq. 55), at finite values of $\kappa$. In Fig. 6 we show the evolution of the boundary order parameter as a function of $\beta$, for $\kappa=2.0$ in (a) and $\kappa=1.0$ in (b), for different system sizes. These reveal a transition from a disordered, symmetric boundary on the confined side (small $\beta$) to an ordered, symmetry-broken boundary on the Higgs side (large $\beta$). This value of $\kappa$ is quite close to the bulk critical point ($\kappa_{c}\sim 0.7$), demonstrating that the boundary phase transition persists far into the phase diagram where the bulk is quite strongly fluctuating. In Fig. 6(c) and (d), we show the Binder cumulant for the same cuts, showing crossing behavior at a $\kappa$-dependent critical coupling. In (e) and (f) we have rescaled $\beta-\beta_{c}(\kappa)$ using the 3D O(4) critical exponent $\nu\approx 0.748$ [37], demonstrating a clean scaling collapse, thus verifying that the transition remains second order and in the same universality class even for relatively small values of $\kappa$. The transition line appears to terminate at the bulk critical point, which can be seen in the red line in Fig. 7, though verifying this numerically is difficult as the bulk correlation length grows larger than the finite width of the open boundaries as the system approaches bulk criticality.
[…]
III.2.3 Tuning the Boundary Coupling
@@
We now consider varying the parameter $\cramped{\alpha=\beta_{\text{bdry}}/\beta_{\text{bulk}}}$, which shifts the location of the $\cramped{\kappa\to\infty}$ transition. Figure 7 shows the behavior of the boundary order parameter and its susceptibility for different values of $\alpha$, along different constant-$\kappa$ cuts at fixed system size. The corresponding behavior of the bulk link susceptibility is shown in black in the background for reference. The bulk transition line moves as $\alpha$ is varied, but appears to remain second-order throughout. Figure 8 summarizes the results by showing the approximate location of the boundary transition line for different values of $\alpha$, with results very similar to the Abelian case (Fig. 4). For $\cramped{\alpha>1}$ the transition moves to smaller values of $\beta_{\text{bulk}}$ and appears to terminate at the bulk critical endpoint (cyan box). No boundary transition is detected for small values of $\kappa$ below the bulk first-order line. For $\cramped{\alpha<1}$ the location of the boundary transition line moves to larger values of $\beta$, and appears to terminates on the line of bulk first-order transitions, at least for sufficiently small $\alpha$.
@@
III.3 Boundary Symmetry Breaking in Non-Abelian Group-Valued Higgs Models
@@
The boundary action is given by Eq. 42. In addition to the global color charge symmetry acting on the boundary, Eq. 44, it also has a global $\mathcal{G}$ symmetry given by right multiplication of the Higgs field
@@
The global symmetry group is therefore $\mathcal{G}_{\text{bdry}}\times\mathcal{G}_{\text{bulk}}$. We expect the bulk phase diagram to be qualitatively similar to the SU(2) case, Fig. 5, with a single thermodynamic phase, a first-order line terminating at a critical endpoint. These models were considered by Fradkin and Shenker [12], who showed that the Higgs and confining regimes are contiguous, as in the Abelian models.
We now consider taking the $\cramped{\kappa\to\infty}$ limit. Maximizing the trace in Eq. 56 yields the constraint $\cramped{\varphi_{i}^{-1}U_{ij}\varphi_{j}=\mathbbm{1}}$, which implies that the bulk links can be expressed in terms of the matter field as
@@
The bulk of the system is completely frozen in this limit, which can be most easily seen in unitary gauge where $\varphi_{i}=\mathbbm{1}$. Substituting Eq. 58 into Eq. 41, the boundary action can then be expressed in terms of the gauge-invariant observables
@@
which is a lattice chiral model. Compare this to the equivalent results for the Abelian case, Eq. 22 and Eq. 51, to which it reduces when $\mathcal{G}=\mathrm{U}(1)$. The $\mathcal{G}_{\text{bdry}}\times\mathcal{G}_{\text{bulk}}$ symmetry acts on this chiral model by left and right multiplication of the $\Theta_{i}$, respectively. Such a model is known to exhibit symmetry breaking to the diagonal $\mathcal{G}$ subgroup [39]. We discuss boundary phase transitions in these models further in Section III.5
Note that if $\mathcal{G}$ is Abelian there is no distinction between left and right group multiplication, therefore $\mathcal{G}_{\text{bulk}}$ and $\mathcal{G}_{\text{bdry}}$ are not independent symmetries of the system. $\mathcal{G}_{\text{bulk}}$ corresponds to global rotation of the Higgs field generated by the total matter charge, Eq. 15, while $\mathcal{G}_{\text{bdry}}$ is generated by the net electric flux through the boundary links, Eq. 17. These two generators are the same operator on the physical gauge-invariant Hilbert space by the Gauss law. In contrast, in the group-valued non-Abelian Higgs models these are really distinct symmetries generated by different physical operators.
III.4 Boundary Symmetry Breaking in Non-Abelian Fundamental-Higgs Models
We now consider non-Abelian gauge groups SU($N$) (and by trivial generalization SO($N$)) with fundamental Higgs fields. As a convenient shorthand, in this section we will represent the Higgs vectors as kets, $\phi_{i}\equiv|\phi_{i}\rangle$, which should not be confused with quantum states. Furthermore, we suppress the subscript “f” on the fundamental representation matrices $U_{\ell}$. The action we write as

### v2

@@ 3.1.1 Higgs Fields
where we take the trace in the fundamental representation. This is a lattice regularization of a gauged principal chiral model [38], a non-linear $\sigma$-model whose target space is the group manifold.
3.1.2 Hamiltonian Formulation and Gauss Law
@@
and $|\hat{E}_{\ell}|^{2}$ and $|\hat{Q}^{\text{m}}_{i}|^{2}$ are the corresponding quadratic Casimir operators, which do not depend on whether we use left- or right-generators. The operator $\hat{W}_{p}$ is the operator analog of Eq.˜27, the trace of the oriented product of $\hat{U}_{\ell}$ on the links of spatial plaquette $p$. Similarly, $\hat{\Lambda}_{\ell}$ is the operator analog of Eq.˜29.
The eigenstates of $|\hat{E}_{\ell}|^{2}$ correspond to the irreducible representations of the gauge group, with the $\hat{U}_{\ell}$ acting as raising and lowering operators [39, 40, 41]. The same is true for the group-valued Higgs, with $\hat{\varphi}_{i}$ acting as the raising and lowering operators, while for the fundamental vector-valued Higgs, the charge eigenstates are angular momentum eigenstates of a rigid rotor.
@@
Note that this is similar but subtly distinct from the Abelian case, Eqs.˜11, 12 and 13, where we used $\hat{E}_{\ell}^{\text{Abelian}}=-\hat{E}^{\text{Abelian}}_{-\ell}$. This relationship is not true in non-Abelian gauge theory. Instead, left- and right-translations of the gauge field are related by
@@
which are manifestly orientation-independent. The classic (though heuristic) way to think of this is that the gauge bosons (gluons) carry a distinct charge and anti-charge in the two directions along the link. In the Abelian case, $U_{\ell}^{\text{a}}=1$ in Eq.˜39, and the link charge Eq.˜40 is exactly zero.
@@
3.1.3 Open Boundaries and Global Symmetry
We introduce electric open boundary conditions as in Fig.˜2, with dynamical links extending from the bulk to the vacuum which allow electric flux to pass through the boundary. We denote the link variables on the boundary link touching site $i$ by $U_{i}$, with the convention that the link is oriented “in” from the vacuum to site $i$. We have minimal open Wilson strings going around the boundary plaquettes, which we can write as
@@ 3.1.3 Open Boundaries and Global Symmetry
Notice that each boundary link only receives a transformation from the inside end, where it terminates on a matter field. In addition to this gauge symmetry, the boundary action Eq.˜42 has a physical global $\mathcal{G}$ symmetry acting on the boundary,
@@
which rotates all matter fields in the fundamental representation and all gauge fields, including boundary links, in the adjoint representation. But this is not the generator of global gauge transformations, Eq.˜43, under which the boundary links only rotate from the inside. The generator of global gauge transformations is the “bulk charge”
@@
where we have assumed global gauge invariance to identify $\hat{Q}^{a}_{\text{bulk}}=0$. Thus the boundary symmetry Eq.˜44 may be viewed, by the Gauss law constraint, as being generated by the total color charge of the system. This is analogous to the Abelian case, where the total charge of the system is just the matter charge, Eq.˜15, since the links do not carry any charge. Note, however, that in the non-Abelian case the charge of the boundary links is “fractionalized” into a piece that contributes to the bulk charge and a piece that contributes to the boundary charge.
[…]
3.2 Boundary Symmetry Breaking in the SU(2) Higgs Model
@@
where $\phi_{1}$ and $\phi_{2}$ are complex numbers. Note that the determinant of $\varphi_{i}^{\text{f}}$ is the length of the rotor. The Wilson link for the vector Higgs model, Eq.˜28, can then be written as
@@
which is exactly equivalent to the group-valued Higgs definition, Eq.˜29. This makes the SU(2) Higgs rotors special, since they may be viewed as either vector-valued or group-valued (which is also the case for U(1) rotor).
3.2.1 Bulk Phase Diagram
To map the phase diagram, we perform classical Monte Carlo simulations for the 4D $\mathrm{SU}(2)$ model defined by the Fradkin-Shenker action, Eq.˜7, with periodic boundary conditions on an $L^{4}$ hypercubic lattice [30, 44, 45]. The phase diagram can be mapped out by measuring the Wilson plaquette and link observables along with their variances (i.e. susceptibilities), which are shown in Fig.˜5. There is a roughly horizontal first-order transition line extending from a critical endpoint at $(\beta_{c},\kappa_{c})\approx(2.73,0.70)$ [42] (cyan box) towards $\cramped{\beta\to\infty}$. This line is clearly visible in the link susceptibility, Fig.˜5(b).
The phase diagram exhibits only one clearly distinct bulk phase—the confined-Higgs phase—a deconfined phase being absent in non-Abelian theories. The phase diagram is, however, roughly separated into two regions indicated by the behavior of the Wilson link expectation value, shown in Fig.˜5(a), with $\cramped{\langle\mathrm{Re}\,\Lambda_{\ell}\rangle\sim 0}$ indicating the confining regime and $\cramped{\langle\mathrm{Re}\,\Lambda_{\ell}\rangle\sim 1}$ indicating the Higgs regime. To the left of this critical endpoint is a supercritical region (a Widom line [46]) extending to smaller $\beta$ and larger $\kappa$, a rapid crossover from the Higgs to the confined regimes. The location of this supercritical region is most evident in the Wilson plaquette susceptibility, Fig.˜5(d), which shows a pronounced intensity emanating from the critical endpoint. We expect this phase diagram to be qualitatively consistent with those for general non-Abelian Higgs models with either fundamental vector- or group-valued Higgs fields. In the group-valued case the continuity of the Higgs and confined regimes was proven by Fradkin and Shenker [14].
@@ 3.2.1 Bulk Phase Diagram
3.2.2 Boundary Symmetry Breaking
We now consider open boundaries, with boundary action Eq.˜42 in a slab geometry. As in Section˜2.3, we start by considering the limiting behavior when $\kappa\to\infty$. In this limit every bulk link satisfies the constraint $\mathrm{Re}\,\Lambda_{\ell}=1$. In this section we will resolve this constraint by fixing a gauge. Gauge-invariant formulations for the group-valued representation are presented in Section˜3.3, and for the fundamental representation in Section˜3.4.
@@ 3.2.2 Boundary Symmetry Breaking
$S^{\text{bdry}}_{\kappa\to\infty}=-\beta\sum_{\langle ij\rangle\in\partial X}S^{\mu}_{i}S^{\mu}_{j}\quad\text{(unitary gauge)}.$ (53)
@@
where the former is an $\mathrm{SU}(2)$ matrix which decomposes according to Eq.˜52, and the latter is explicitly a 4-component unit-length vector. For large but finite $\kappa$, the bulk fluctuations are strongly gapped and the boundary should behave as quasi-(D1)(D-1)-dimensional, and we expect the boundary phase transition to persist as in the Abelian case.
@@
To test this prediction, we perform Monte Carlo simulations with open boundaries and measure the order parameter, Eq.˜54 (defined in terms of gauge invariant observables Eq.˜55), at finite values of $\kappa$. In Fig.˜6 we show the evolution of the boundary order parameter as a function of $\beta$, for $\kappa=2.0$ in (a) and $\kappa=1.0$ in (b), for different system sizes. These reveal a transition from a disordered, symmetric boundary on the confined side (small $\beta$) to an ordered, symmetry-broken boundary on the Higgs side (large $\beta$). This value of $\kappa$ is quite close to the bulk critical point ($\kappa_{c}\sim 0.7$), demonstrating that the boundary phase transition persists far into the phase diagram where the bulk is quite strongly fluctuating. In Fig.˜6(c) and (d), we show the Binder cumulant for the same cuts, showing crossing behavior at a $\kappa$-dependent critical coupling. In (e) and (f) we have rescaled $\beta-\beta_{c}(\kappa)$ using the 3D O(4) critical exponent $\nu\approx 0.748$ [43], demonstrating a clean scaling collapse, thus verifying that the transition remains second order and in the same universality class even for relatively small values of $\kappa$. The transition line appears to terminate at the bulk critical point, which can be seen in the red line in Fig.˜7, though verifying this numerically is difficult as the bulk correlation length grows larger than the finite width of the open boundaries as the system approaches bulk criticality.
3.2.3 Tuning the Boundary Coupling
@@
We now consider varying the parameter $\cramped{\alpha=\beta_{\text{bdry}}/\beta_{\text{bulk}}}$, which shifts the location of the $\cramped{\kappa\to\infty}$ transition. Figure˜7 shows the behavior of the boundary order parameter and its susceptibility for different values of $\alpha$, along different constant-$\kappa$ cuts at fixed system size. The corresponding behavior of the bulk link susceptibility is shown in black in the background for reference. The bulk transition line moves as $\alpha$ is varied, but appears to remain second-order throughout. Figure˜8 summarizes the results by showing the approximate location of the boundary transition line for different values of $\alpha$, with results very similar to the Abelian case (Fig.˜4). For $\cramped{\alpha>1}$ the transition moves to smaller values of $\beta_{\text{bulk}}$ and appears to terminate at the bulk critical endpoint (cyan box). No boundary transition is detected for small values of $\kappa$ below the bulk first-order line. For $\cramped{\alpha<1}$ the location of the boundary transition line moves to larger values of $\beta$, and appears to terminates on the line of bulk first-order transitions, at least for sufficiently small $\alpha$.
Figure 8: Location of the boundary phase transition for the SU(2) Higgs model for different values of $\alpha=\beta_{\text{bdry}}/\beta_{\text{bulk}}$, determined by the location of the peak in the boundary order parameter susceptibility (cf. Fig.˜7) . Large values of $\alpha$ push the transition line to smaller $\beta_{\text{bulk}}$, while small values of $\alpha$ push the phase boundary to larger $\beta_{\text{bulk}}$. The cyan square indicates the location of the bulk critical endpoint. For $\alpha=1$ (red), the boundary transition line appears to terminate at the bulk critical point, and closely follows the bulk rapid-crossover (super critical) region that extends beyond the critical endpoint. This remains true for $\alpha>1$, while for sufficiently small $\alpha<1$ the boundary transition line appears to terminate on the bulk first-order transition line.
@@
3.3 Boundary Symmetry Breaking in Non-Abelian Group-Valued Higgs Models
@@
The boundary action is given by Eq.˜42. In addition to the global color charge symmetry acting on the boundary, Eq.˜44, it also has a global $\mathcal{G}$ symmetry given by right multiplication of the Higgs field
@@
The global symmetry group is therefore $\mathcal{G}_{\text{bdry}}\times\mathcal{G}_{\text{bulk}}$. We expect the bulk phase diagram to be qualitatively similar to the SU(2) case, Fig.˜5, with a single thermodynamic phase, a first-order line terminating at a critical endpoint. These models were considered by Fradkin and Shenker [14], who showed that the Higgs and confining regimes are contiguous, as in the Abelian models.
We now consider taking the $\cramped{\kappa\to\infty}$ limit. Maximizing the trace in Eq.˜56 yields the constraint $\cramped{\varphi_{i}^{-1}U_{ij}\varphi_{j}=\mathbbm{1}}$, which implies that the bulk links can be expressed in terms of the matter field as
@@
The bulk of the system is completely frozen in this limit, which can be most easily seen in unitary gauge where $\varphi_{i}=\mathbbm{1}$. Substituting Eq.˜58 into Eq.˜41, the boundary action can then be expressed in terms of the gauge-invariant observables
@@
which is a lattice chiral model. Compare this to the equivalent results for the Abelian case, Eq.˜22 and Eq.˜51, to which it reduces when $\mathcal{G}=\mathrm{U}(1)$. The $\mathcal{G}_{\text{bdry}}\times\mathcal{G}_{\text{bulk}}$ symmetry acts on this chiral model by left and right multiplication of the $\Theta_{i}$, respectively. Such a model is known to exhibit chiral symmetry breaking, i.e. breaking to the diagonal $\mathcal{G}$ subgroup [47]. We discuss boundary phase transitions in these models further in Section˜3.5
Note that if $\mathcal{G}$ is Abelian there is no distinction between left and right group multiplication, therefore $\mathcal{G}_{\text{bulk}}$ and $\mathcal{G}_{\text{bdry}}$ are not independent symmetries of the system. $\mathcal{G}_{\text{bulk}}$ corresponds to global rotation of the Higgs field generated by the total matter charge, Eq.˜15, while $\mathcal{G}_{\text{bdry}}$ is generated by the net electric flux through the boundary links, Eq.˜17. These two generators are the same operator on the physical gauge-invariant Hilbert space by the Gauss law. In contrast, in the group-valued non-Abelian Higgs models these are really distinct symmetries generated by different physical operators.
3.4 Boundary Symmetry Breaking in Non-Abelian Fundamental-Higgs Models
We now consider non-Abelian gauge groups $\mathrm{SU}(N)$ (and by trivial generalization $\mathrm{SO}(N)$) with fundamental Higgs fields. As a convenient shorthand, in this section we will represent the Higgs vectors as kets, $\phi_{i}\equiv|\phi_{i}\rangle$, which should not be confused with quantum states. Furthermore, we suppress the subscript “f” on the fundamental representation matrices $U_{\ell}$. The action we write as